OpenStudy (anonymous):

Which expression is equivalent to 3(m − 3) + 4?

OpenStudy (alexandervonhumboldt2):

use distibutive property first: a*(b-c)=ab-ac

OpenStudy (alexandervonhumboldt2):

@HeartIce

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

How

OpenStudy (anonymous):

You don't know it? you must be new to distributing

OpenStudy (alexandervonhumboldt2):

3m-3*3=3m-9 now add 4 to it 3m-9+4 what you get?

OpenStudy (anonymous):

Not really i would do it but i dont really undersand it as much as i use to

OpenStudy (anonymous):

2.25?

OpenStudy (alexandervonhumboldt2):

-9+4=-5

OpenStudy (alexandervonhumboldt2):

3m-5

OpenStudy (anonymous):

oh thank u a million

OpenStudy (anonymous):

Is best response medal?

OpenStudy (anonymous):

During his workouts last week, Allen jogged the same amount of time each day for 6 days. Use the following information to determine the number of minutes Allen jogged each day: During each workout, he jogged and he rode his bike. He rode his bike for 45 minutes each day. He worked out for a total of 360 minutes last week. What is the number of minutes that Allen jogged each day last week? Enter your answer in the box below.

OpenStudy (anonymous):

PLZ HELP...AGAIN >.> <.<

OpenStudy (anonymous):

3m -3+4

OpenStudy (anonymous):

oh y thank u

OpenStudy (anonymous):

have a nice day

OpenStudy (anonymous):

oh wait i have more questions

OpenStudy (anonymous):

Marla bought one necklace and two bracelets for $67.50. Use the following information to determine the amount she paid for each bracelet: The price of the necklace was $27. The bracelets were both the same price. Enter your numerical answer only in the box; do not include the dollar sign.

OpenStudy (anonymous):

@Lakeekee_

OpenStudy (anonymous):

@Frostbite

OpenStudy (anonymous):

which question do you need help on

OpenStudy (anonymous):

The Marla one plz

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

do you wanna no how to do it or just the answer

OpenStudy (anonymous):

Take your time

OpenStudy (anonymous):

anything thats best with u

OpenStudy (anonymous):

Im very bad at math lol

OpenStudy (anonymous):

 is this what im suppose to do?